OpenStudy (anonymous):

Identify the slope and y-intercept: of the inequality 8x + 2y > 12. A) slope = 4, y-intercept: = 6 B) slope = -4, y-intercept: = 12 C) slope = -4, y-intercept: = 6 D) slope = -1/4 , y-intercept: = 12

OpenStudy (anonymous):

Ok... u try to solve for y 8x + 2y > 12 2y > -8x + 12 y = -4x + 6 now, its in the form of y = mx + b! m is your slope, b is your y-intercept! now, do you know whats your slope and whats your y-intercept?

OpenStudy (anonymous):

y intercept is 6 and slope is -4?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

yay!! thanks!

OpenStudy (anonymous):

could you help me with another question?

OpenStudy (anonymous):

ok